Janet has five fewer fifteen-cent stamps than twenty-cent stamps. If the total value of the stamps is $3.45, how many twenty-cent stamps does she have?
Answer by checkley77(12844)

.15x+.20y=3.45
x=y-5
.15(y-5)+.20y=3.45
.15y-.75+.20y=3.45
.35y=3.45+.75
.35y=4.20
y=4.20/.35
y=12 20 cent stamps.
x=12-5=7 15 cent stamps
Proof:
.15*7+.20*12=3.45
1.05+2.40=3.45
3.45=3.45